ABC currently pays no dividends, choosing instead to re-invest all earnings in the firm. However, the firm anticipates that beginning in year 6, they will run out of profitable investments and begin paying a dividend of $5.00 per share. They anticipate that dividends will grow at 2% per year in perpetuity following that.

The required return on the cash flows of ABC’s dividends is 15% per year.

  1. What is the equilibrium value of a share of ABC today?
  2. Assuming that nothing about ABC’s cash flows, or its required return, changes, what should be the equilibrium value of a share of ABC at t=5? At t=6?
Answer #1

1.
=5/(15%-2%)*1/1.15^5=19.12218213

2.
At t=5: =5/(15%-2%)=38.46153846

At t=6: =5*(1+2%)/(15%-2%)=39.23076923